OpenStudy (thickgirl_angel):

A radio active isotope decays according to the exponential decay equation where t is in days. Round to the thousandths place. For the half life: The half life is the solution (t) of the equation : a2=ae−9.548ta

OpenStudy (kgrendel0324):

\[a/2=ae ^{-3.524t} \] \[1/2=e ^{-3.524t}\]

OpenStudy (kgrendel0324):

ln(0.5) = -3.5324t t = -ln(0.5)/3.5324 t =~ 0.1962 days

OpenStudy (kgrendel0324):

I solved this for you yesterday. Did you need help?